OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. A = 80°, a = 24, b = 50

OpenStudy (anonymous):

Zero

OpenStudy (anonymous):

how ?

OpenStudy (callmekiki):

no direct answers please.

OpenStudy (callmekiki):

80 + 24 + 50 = 154 All the angles in a triangle have to equal 180.

OpenStudy (anonymous):

80+24= 104 104+50=154

OpenStudy (anonymous):



OpenStudy (anonymous):

you can also use sin formula \[\frac{ a }{ \sin A }=\frac{ b }{ \sin B }=\frac{ c }{ \sin C }\]